Find the resultant vector using component method:




A = 20km, Eastward



B= 15km, 35 degrees North of West

1
Expert's answer
2021-11-22T10:13:11-0500

The components:


"A_E=20\text{ km},\\
A_N=0.\\
B_E=-15\cos35°=-12.3\text{ km},\\
B_N=15\sin35°=8.6\text{ km}."

The resultant is


"|r|=\sqrt{(A_E+B_E)^2+(A_N+B_N)^2}=11.54\text{ km},\\\space\\
\theta=\arctan\frac{A_N+B_N}{A_E+B_E}=48.16°\text{ N of E}."